OpenStudy (anonymous):

r = 3secθ to rectangular form???

OpenStudy (anonymous):

x^2+y^2+3y=0, right?

OpenStudy (phi):

I would use x^2 + y^2 = r^2, x= r cos θ , y = r sin θ in this case r = 3 /cos θ replace cos θ = x/r to get r = 3/ (x/r) r = 3r /x cancel r from both sides 1 = 3/x x = 3

OpenStudy (anonymous):

Oh...

OpenStudy (anonymous):

So for rsecθ = 3, is it x=-3?

OpenStudy (phi):

r sec = 3 r/cos = 3 r / (x/r) =3 r^2 / x = 3 (x^2 + y^2)/x = 3 x^2 -3x + y^2 = 0 complete the square on the x terms, to make this look nicer (equation of a circle)